The vector joing the points A(1,1,-1) and B(2,-3,4) and pointing from A to B-

(a) -i+4j-5k
(b) i+4j+5k
(c) i-4j+5k

Grade:11

1 Answers

rupali sehgal
11 years ago

clearly, OA= i+j-k

OB= 2i-3j+4k

by triangle law,

OA+AB=OB

AB=OB-OA

AB= 2i-3j+4k -(i+j-k)= (2i-i) + (-3j-j) + (4k+k)= i-4j+5k

Therefoe, correct option is (C)